### v1

@@ Abstract
Drug repositioning is an economically efficient strategy used to discover new indications for existing drugs beyond their original approvals, expanding their applicability and usage to address challenges in disease treatment. In recent years, deep-learning techniques for drug repositioning have gained much attention. While most deep learning-based research methods focus on encoding drugs and diseases by extracting feature information from neighbors in the network, they often pay little attention to the potential relationships between the features of drugs and diseases, leading to imprecise encoding of drugs and diseases. To address this, we design a dual-feature drug repositioning neural network (DFDRNN) model to achieve precise encoding of drugs and diseases. DFDRNN uses two features to represent drugs and diseases: the similarity feature and the association feature. The model incorporates a self-attention mechanism to design two dual-feature extraction modules for achieving precisely encoding of drugs and diseases: the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module. The IntraDDFE module extracts features from a single domain (drug or disease domain), while the InterDDFE module extracts features from the mixed domain (drug and disease domain). In particular, the feature is changed by InterDDFE, ensuring a precise encoding of drugs and diseases. Finally, a cross-dual-domain decoder is designed to predict drug-disease associations in both the drug and disease domains. Compared to six state-of-the-art methods, DFDRNN outperforms others on four benchmark datasets, with an average AUROC of 0.946 and an average AUPR of 0.597.
@@
Drug repositioning, neural network, dual-feature, association feature, cross-dual-domain decoder
@@ 1 Introduction
In today’s Medical field, drug development is complex and time-consuming. Developing a new drug involves three stages: the discovery stage, the preclinical stage, and the clinical stage, which typically takes 10-15 years, with an average cost of up to \$2.8 billion fisher2006alterations ; vijayan2022enhancing . Although scientists continue to discover new drug molecules through unremitting efforts and innovation, the traditional drug development model often faces many risks, including the termination of drug development due to drug efficacy, drug safety, or commercial concerns. Remarkably, a significant proportion (80-90%) of these projects fail in clinical trials vijayan2022enhancing ; dimasi2010trends . Therefore, finding a more efficient and precise method for drug development has become critically important.
Drug repositioning, a strategy that garnered significant attention, is proposed to explore alternative indications and potential side effects for already-licensed medications dudley2011exploiting ; wang2024landscape . Compared with traditional drug research and development methods, this approach offers several advantages, such as shorter time, lower cost, and reduced risk. Computational drug repositioning, a favored method among researchers, can significantly reduce the time and energy required to identify candidate drugs. Computational drug repositioning methods are divided into four categories cai2023machine ; luo2021biomedical : network diffusion-based methods, machine learning-based methods, matrix factorization-based methods, and deep learning-based methods.
Network diffusion-based methods classify the drug reposition-ing problem as a link prediction issue within graph neural networks kumar2020link . These methods focus on building heterogeneous networks containing multiple types of biological entities and linking drugs to diseases through message propagation on paths bridging different networks. A commonly used method is a random walk and its variants luo2021biomedical ; lagcn . For instance, Martinez et al. martinez2015drugnet proposed a network-based drug repositioning prioritization method called DrugNet. It integrates information about diseases, drugs, and proteins to simultaneously prioritize drug-disease and disease-drug associations. Another example is Luo et al.’s method BIRW_Cdataset , which effectively predicted potential associations between drugs and diseases by constructing an integrated similarity network between drugs and diseases and combining it with a bidirectional random walk (BiRW).
Machine learning-based methods classify drug repositioning as a binary classification problem. These methods typically use traditional machine learning techniques for drug-disease association prediction, including random forest (RF), support vector machine (SVM), and decision trees cai2023machine . For example, Peng et al. screening designed a novel DTI screening framework called PUDTI. PUDTI utilizes SVM to infer candidate targets for drug repositioning. Shi et al. shi2019lasso proposed a method for predicting drug-target interactions based on Lasso dimensionality reduction and a random forest classifier. Zhao et al. HINGRL proposed HINGRL, which uses a random forest classifier in the final decoding step to predict drug-disease association. In addition, there are decision trees gradient , linear regression, and logistic regression you2019predicting ; elisei2019recursive .
Matrix decomposition-based methods categorize the drug repositioning problem as a recommendation system issue. These methods employ matrix factorization techniques to reduce dimensionality and extract key features, thus better capturing the associations between drugs and diseases. Some standard matrix factorization techniques include singular value decomposition (SVD) and non-negative matrix factorization (NMF). For example, Yang et al. MSBMF proposed a Multiple Similarity Bilinear Matrix Factorization (MSBMF) method, where the similarity matrices of drugs and diseases are concatenated and the drug-disease association matrix is reshaped through matrix decomposition. Luo et al. luo2018SVT employed a fast Singular Value Thresholding (SVT) algorithm to propose a drug repositioning recommendation system (DRRS). Additionally, Sadegh et al. sadeghi2022NMF-DR and Xuan et al. xuan2019dd combined non-negative matrix factorization to predict drug-disease associations.
In recent years, methods based on deep learning have become a hot research topic for drug repositioning. They can effectively capture high-level hidden representations from various data sources and demonstrate state-of-the-art performance. Unlike traditional machine learning methods, deep learning approaches can learn more abstract information by building deep architectures without manual feature selection and tuning yu2022deep . Generally, methods based on deep learning for drug repositioning often integrate graph neural networks such as Graph Convolutional Networks (GCN) and Graph Attention Networks (GAT) to construct models. For example, Yu et al. lagcn integrated a drug-disease heterogeneous network with a GCN network to build a multi-layer connected end-to-end network model called LAGCN. They also incorporated layer attention mechanisms for drug-disease association prediction. Gu et al. gu2022REDDA designed a biological network of five biological entities and equipped it with three attention mechanisms to predict drug-disease associations. Wang et al. wang2023MSGCA designed a multi-similarity graph convolutional autoencoder(MSGCA). Sun et al. sun2022PSGCN predicted new drug-disease associations by constructing partner-specific graphs for drug-disease pairs and combining them with the GCN. By incorporating the GCN, Xuan et al. xuan2022heterogeneous constructed multiple drug-disease heterogeneous networks based on different drug similarities and encoded drugs and diseases at Multi-scale neighbors. Cai et al. cai2021DRHGCN designed intra-domain and inter-domain feature extraction modules through the GCN and the bilinear aggregator, proposing DRHGCN. Meng et al. DRAGNN utilized a GAT for encoding drugs and diseases and introduced a weighted local information augmented graph neural network model named DRAGNN. Most current deep learning-based methods for drug repositioning typically utilize the similarity between neighbors and aggregate neighbor features for training, achieving considerable results. However, many end-to-end deep learning models do not involve purposeful encoding. They delegate the task to the optimizer, relying solely on the coordinated training of the encoding and decoding processes to complete association prediction. This may lead to the introduction of a large amount of noisy information during model training, resulting in imprecise encoding of drugs and diseases and affecting the prediction results.
@@
This paper proposes a dual-feature drug repositioning neural network (DFDRNN) model to address the above issue. The DFDRNN constructs two features for drugs and diseases: the similarity feature and the association feature. It utilizes the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module to extract drug and disease features and achieve precise encoding. A self-attention mechanism (SAM) is employed in these two modules to capture the complex relationships between adjacent nodes. Finally, the decoder adopts a cross-dual-domain decoder to achieve precise prediction of drug-disease associations. Compared to GCN-based methods, DFDRNN employs a SAM to dynamically adjust the attention level of each input element to neighbor information, providing fine-grained weight allocation. This enables efficient aggregation of neighbors and enhances the model’s predictive power. We compare DFDRNN with six state-of-the-art methods on four benchmark datasets. Experimental results show that our model outperforms other models. This article makes the following contributions:
(1) We propose a dual-feature method to represent drugs and diseases. The proposed approach involves two features: the similarity feature and the association feature. These two features are independent of each other in the network, while they can be extracted simultaneously to represent drugs and diseases in each network layer.
(2) Two dual-feature extraction modules are proposed: the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module. In particular, the InterDDFE module changes the similarity (or association) feature into the association (or similarity) feature. The proposed two modules effectively capture the features of drugs and diseases, ensuring a precise encoding of drugs and diseases.
@@
(3) A cross-dual-domain decoder is proposed, implementing a cross-computation between two encoded features to obtain decoding results for the drug and disease domains. The fusion of the two decoding results improves the accuracy of association prediction.
@@
The remainder of this paper is organized as follows. Section 2 describes the datasets and methods, detailing the implementation process of DFDRNN; Section 3 clarifies the experimental design and analyzes the experimental results; and Section 4 concludes our work.
@@ 2.1 Datasets
We conducted experiments on four benchmark datasets. The first dataset, Fdataset Fdataset , includes 593 drugs from DrugBank knox2024drugbank and 313 diseases listed in the Online Mendelian Inheritance in Man (OMIM) hamosh2005OMIM database, comprising 1933 drug-disease associations. The second dataset, Cdataset BIRW_Cdataset , consists of 663 drugs registered in DrugBank, 409 diseases from the OMIM database, and 2532 known drug-disease associations. The third dataset, Ldataset lagcn , contains 18,416 drug-disease associations between 269 drugs and 598 diseases sourced from the Comparative Toxicogenomics Database(CTD) davis2023CTD . The fourth dataset, LRSSL liang2017lrssl , comprises 763 drugs, 681 diseases, and 3051 validated drug-disease associations. Table 1 summarizes the four datasets’ statistics.
This study determines drug similarity based on chemical structures obtained from DrugBank using standardized SMILES notation weininger1988smiles . The Tanimoto score of the 2D chemical fingerprints of two drugs is then calculated using the chemistry development kit steinbeck2003chemistry . Disease similarity, on the other hand, is based on disease phenotypes and is computed using MimMiner van2006text . MimMiner gauges disease similarity by analyzing the similarity of MeSH terms in medical descriptions of diseases from the OMIM database lipscomb2000medical .
@@ 2.2 Model structure
Our research is focused on predicting potential associations between drugs and diseases. To this end, we proposed a dual-feature drug repositioning neural network (DFDRNN) model. The encoder of DFDRNN encompasses two dual-feature extraction modules: the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module. While the IntraDDFE module extracts features among drugs (or diseases), the InterDDFE module extracts features between drugs and diseases and changes the feature. A self-attention (SAM) model is employed when aggregating information to capture the complex relationship among adjacent nodes. The decoder of DFDRNN receives the results from the encoder and implements cross-dual-domain decoding, including both the drug and disease domains. Figure 1 illustrates the diagram of the proposed DFDRNN model.
Figure 1: DFDRNN generates the similarity feature and the association feature for drugs and diseases based on established similarity networks and association networks. The encoding process utilizes a multi-layer connected network architecture, with each layer performing IntraDDFE and InterDDFE based on self-attention. A layer attention mechanism is also introduced to compute the final similarity and association features. Finally, the decoder calculates drug-disease association scores by decoding separately for the drug and disease domains.
@@ 2.2.1 Three networks
To illustrate the relationships among biological entities (drugs and diseases) and describe their feature, we establish three networks: the drug-drug similarity network, a weighted complete graph with a node-set representing the drugs, and the weight of each edge representing the degree of similarity of the two drug incident with the edge; the disease-disease similarity network, a weighted complete graph with node set representing the diseases and the weight of each edge representing the degree of similarity of the two diseases incident with the edge; and the drug-disease association network, a bipartite network with two parts representing the sets of drugs and diseases, respectively and each edge representing the associativity between the drug and diseases incident with the edge. We use three matrices to represent these three networks.
Specifically, suppose that there are $n$ drugs (denoted by $r_{1},r_{2},\ldots,r_{n}$) and $m$ diseases (denoted by $d_{1},d_{2},\ldots,d_{m}$). We use $S^{r}$ (or $S^{d}$) to denote the adjacency matrix of a drug-drug similarity network (or disease-disease similarity network), which is a $n\times n$ ($m\times m$) real matrix such that the entry $s_{ij}\in[0,1]$ denotes the degree of similarity of drug $i$ and drug $j$ (or disease $i$ and disease $j$), where $i,j\in\{1,2,\ldots,n\}$ (or $i,j\in\{1,2,\ldots,m\}$). For a drug-disease association network, we use a $n\times m$ binary matrix, denoted by $A$, to represent it, where the entry $a_{ij}=1$ if drug $i$ is associated with disease $j$ and $a_{ij}=0$ if drug $i$ is not associated with disease $j$ or the associativity is yet to be observed, for $i=1,2,\ldots,n$ and $j=1,2,\ldots,m$.
To reduce the impact of data noise, we aggregate only the top-$t$ neighbors for each drug in $S^{r}$ and each disease in $S^{d}$, and then two new matrices $R\in\mathbb{R}^{n\times n}$ and $D\in\mathbb{R}^{m\times m}$ are generated to replace $S^{r}$ and $S^{d}$, respectively, where the entries $r_{ij}$ ($i,j\in\{1,2,\ldots,n\}$) of $R$ and $d_{ij}$ ($i,j\in\{1,2,\ldots,m\}$) of $D$ are defined as follows. For a drug $r_{i}$ (and a disease $d_{i}$), we denote by $N_{k}(r_{i})$ (and $N_{k}(d_{i})$) the set of the $t$-nearest neighbors of $r_{i}$ in $S^{r}$ (and $d_{i}$ in $S^{d}$), i.e., the weight of edge $r_{i}x$ (and edge $d_{i}y$) for every $x\in N_{k}(r_{i})$ (and every $y\in N_{k}(d_{i})$) is top-$t$ among all neighbors of $r_{i}$ (and $d_{i}$), and let $\tilde{N_{k}}(r_{i})=N_{k}(r_{i})\cup\{r_{i}\}$ (and $\tilde{N_{k}}(d_{i})=N_{k}(d_{i})\cup\{d_{i}\}$) be the extended $t$-nearest neighbors of $r_{i}$ (and $d_{i}$).
$\begin{matrix}R_{ij}=\begin{cases}1,&\mbox{if }r_{j}\in{\tilde{N_{k}}(r_{i})}\\ 0,&\text{otherwise}\end{cases}\quad&D_{ij}=\begin{cases}1,&\mbox{if }d_{j}\in{\tilde{N_{k}}(d_{i})}\\ 0,&\text{otherwise}\end{cases}\end{matrix}$ (1)

### v2

@@ Abstract
Uncovering new therapeutic uses of existing drugs, drug repositioning offers a fast and cost-effective strategy and holds considerable significance in the realm of drug discovery and development. In recent years, deep learning techniques have emerged as powerful tools in drug repositioning due to their ability to analyze large and complex datasets. However, many existing methods focus on extracting feature information from nearby nodes in the network to represent drugs and diseases, without considering the potential inter-relationships between the features of drugs and diseases, which may lead to inaccurate representations. To address this limitation, we use two features (similarity and association) to capture the potential relationships between the features of drugs and diseases, proposing a Dual-Feature Drug Repositioning Neural Network (DFDRNN) model. DFDRNN uses a self-attention mechanism to extract neighbor features and incorporates two dual-feature extraction modules: the intra-domain dual-feature extraction (IntraDDFE) module for extracting features within a single domain (drugs or diseases) and the inter-domain dual-feature extraction (InterDDFE) module for extracting features across domains. By utilizing these modules, we ensure more appropriate encoding of drugs and diseases. Additionally, a cross-dual-domain decoder is designed to predict drug-disease associations in both domains. Our proposed DFDRNN model outperforms six state-of-the-art methods on four benchmark datasets, achieving an average AUROC of 0.946 and an average AUPR of 0.597. Case studies on two diseases show that the proposed DFDRNN model can be applied in real-world scenarios, demonstrating its significant potential in drug repositioning.
@@
[…]
@@ 1 Introduction
In the field of medicine, drug development is a complex and time-consuming process. It involves three main stages: the discovery stage, the preclinical stage, and the clinical stage. This process typically takes 10-15 years and around \$2.8 billion on average (Fisher Wilson, 2006; Vijayan et al., 2022). Despite continuous efforts and innovation, the traditional drug development model faces significant risks and sometimes leading to the termination of projects due to concerns about drug efficacy, safety, or commercial viability. Notably, a high proportion (80-90%) of these projects fail during clinical trials (Vijayan et al., 2022; DiMasi et al., 2010). Therefore, finding a more efficient and precise method for drug development is critically important and valuable.
Computational drug repositioning is a powerful strategy that aims to discover alternative uses of already approved drugs taking into account their potential side effects (Dudley et al., 2011; Wang et al., 2024). This approach uses computational and bioinformatics tools and is more efficient and cost-effective than traditional drug research and development methods. It allows for rapid screening of numerous candidate drugs and accelerates the development of new therapeutic options. Repurposing existing drugs can significantly reduce development costs compared to developing new drugs from scratch. In response to new health crises, such as pandemics, repurposing known drugs can provide a quicker response compared to developing new treatments. This method can also identify new usage for drugs that may be effective for diseases lacking effective treatments and thereby addressing significant health challenges.
Machine learning is commonly used in computational drug repositioning. Traditional techniques like random forest (RF), support vector machine (SVM), and decision trees are often applied for this purpose (Cai et al., 2023). For instance, Peng et al. (2017) introduced a novel drug-target interaction screening framework called PUDTI, which utilized SVM to identify potential drug repositioning targets. Shi et al. (2019) proposed a method for predicting drug-target interactions using Lasso dimensionality reduction and a random forest classifier. Zhao et al. (2022) and Yang et al. (2023) employed random forests in their final prediction step, while other techniques such as decision trees (Xuan et al., 2019), matrix factorization(Yang et al., 2021; Luo et al., 2018; Sadeghi et al., 2022), and logistic regression (Yu et al., 2020) have also been developed.
In recent years, deep learning technology has become a powerful tool in bioinformatics, as evidenced by (Zhang et al., 2017; Mahdaddi et al., 2021; Bansal et al., 2024). Unlike traditional machine learning methods, deep learning technology can effectively capture high-level hidden representation using deep architectures without manual feature selection and tuning (Yu et al., 2022). Existing deep learning methods for drug repositioning often incorporate graph neural networks like Graph Convolutional Networks (GCN) and Graph Attention Networks (GAT) to for model building (Yu et al., 2021; Cai et al., 2021; Sun et al., 2022; Tang et al., 2023; Zhang et al., 2024). For instance, Yu et al. (2021) performed graph convolution operations on the drug-disease heterogeneous network to encode drugs and diseases, but did not account for the differences in encoding from different network spaces. Cai et al. (2021) distinguished between intra-domain and inter-domain feature extraction on the drug-disease heterogeneous network for encoding drugs and diseases. Zhang et al. (2024) extracted features of drugs and diseases from multiple neighborhood spaces and designed a contrastive learning loss to fuse these features. However, these studies did not consider the calculation of information weights from different neighbors. Authors in (Gu et al., 2022; Meng et al., 2024; Huang et al., 2024) used graph attention networks (GAT) to calculate the weights of neighbor aggregation and extract key neighbor information for encoding. Most deep learning models have so far been focused on extracting valuable features from neighborhoods. However, these models often employ single encoding approach and lack intentional encoding and decoding processes for association prediction. This can lead to the introduction of a large amount of noisy information during model training, resulting in the imprecise encoding of drugs and diseases, ultimately affecting the accuracy of prediction results.
@@
In order to address this challenge, we propose a Dual-Feature Drug Repositioning Neural Network (DFDRNN) model that uses two features, similarity, and association, to accurately represent drugs and diseases. In comparison to existing methods, the proposed model incorporates two dual-feature extraction modules: the IntraDDFE module for extracting features within a single domain (drug or disease) and the InterDDFE module for extracting features across domains. DFDRNN uses a self-attention mechanism (SAM) to dynamically adjust the attention level of each input element to neighbor information, providing fine-grained weight allocation. This allows for efficient aggregation of neighbors and enhances the model’s predictive power. Additionally, a cross-dual-domain decoder is designed to predict drug-disease associations in both domains. The key contributions of this paper are:
(1) To predict potential associations between drugs and diseases, we propose a dual-feature drug repositioning neural network (DFDRNN) model. A self-attention mechanism is used capture complex relationship among adjacent nodes. The encoder of DFDRNN consists of two dual-feature extraction modules: the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module.
(2) The IntraDDFE module and InterDDFE module extract features exploiting information within a single domain and across domains, respectively, and track the dynamic changes of dual-feature to realize precise encoding of drugs and diseases.
@@
(3) A cross-dual-domain decoder is proposed to perform cross-computation between two encoded features. This process yields decoding results for both the drug and disease domains, which help effectively predicting the drug-disease associations.
(4) The performance of the proposed method is compared with that of six state-of-the-art methods on four benchmark datasets. The experimental results show that DFDRNN outperforms other models. In addition, case studies on Alzheimer’s and Parkinson’s diseases validated the reliability of DFDRNN in practical applications.
@@
The remainder of this paper is organized as follows. Section 2 describes the datasets and methods, detailing the implementation process of DFDRNN; Section 3 clarifies the experimental design and analyzes the experimental results, and Section 4 concludes this work with future research directions.
@@ 2.1 Datasets
We conducted experiments on four benchmark datasets. The first dataset, Fdataset (Gottlieb et al., 2011), includes 593 drugs from DrugBank (Knox et al., 2024) and 313 diseases listed in the Online Mendelian Inheritance in Man (OMIM) (Hamosh et al., 2005) database, comprising 1933 drug-disease associations. The second dataset, Cdataset (Luo et al., 2016), consists of 663 drugs registered in DrugBank, 409 diseases from the OMIM database, and 2532 known drug-disease associations. The third dataset, Ldataset (Yu et al., 2021), contains 18,416 drug-disease associations between 269 drugs and 598 diseases sourced from the Comparative Toxicogenomics Database(CTD) (Davis et al., 2023). The fourth dataset, LRSSL (Liang et al., 2017), comprises 763 drugs, 681 diseases, and 3051 validated drug-disease associations. Table 1 summarizes the four datasets’ statistics.
This study determines drug similarity based on chemical structures obtained from DrugBank using standardized SMILES notation (Weininger, 1988). The Tanimoto score of the 2D chemical fingerprints of two drugs is then calculated using the chemistry development kit (Steinbeck et al., 2003). Disease similarity, on the other hand, is based on disease phenotypes and is computed using MimMiner (Van Driel et al., 2006). MimMiner gauges disease similarity by analyzing the similarity of MeSH terms in medical descriptions of diseases from the OMIM database (Lipscomb, 2000).
@@ 2.2 Model structure
Our research is focused on predicting potential associations between drugs and diseases. To this end, we proposed a dual-feature drug repositioning neural network (DFDRNN) model. The encoder of DFDRNN encompasses two dual-feature extraction modules: the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module. While the IntraDDFE module extracts features among drugs (or diseases), the InterDDFE module extracts features between drugs and diseases and changes the features. A self-attention mechanism (SAM) model is employed when aggregating information to capture the complex relationship among adjacent nodes. The decoder of DFDRNN receives the results from the encoder and implements cross-dual-domain decoding, including both the drug and disease domains. Figure 1 illustrates the diagram of the proposed DFDRNN model.
Figure 1: DFDRNN generates the similarity feature and the association feature for drugs and diseases based on established similarity networks and association networks. The encoding process utilizes a multi-layer connected network architecture, with each layer performing IntraDDFE and InterDDFE based on the self-attention mechanism. A layer attention mechanism is also introduced to compute the final similarity and association features. Finally, the decoder calculates drug-disease association scores by decoding separately for the drug and disease domains.
@@ 2.2.1 Three networks
To illustrate the relationships among biological entities (drugs and diseases) and describe their feature, we establish three networks: the drug-drug similarity network, a weighted complete graph with a node-set representing the drugs, and the weight of each edge representing the degree of similarity between the two drugs incident with the edge; the disease-disease similarity network, a weighted complete graph with node set representing the diseases and the weight of each edge representing the degree of similarity of the two diseases incident with the edge; and the drug-disease association network, a bipartite network with two parts representing the sets of drugs and diseases, respectively and each edge representing the associativity between the drug and diseases incident with the edge. We use three matrices to represent these three networks.
Specifically, suppose that there are $n$ drugs (denoted by $r_{1},r_{2},\ldots,r_{n}$) and $m$ diseases (denoted by $d_{1},d_{2},\ldots,d_{m}$). We use $S^{r}$ (or $S^{d}$) to denote the adjacency matrix of a drug-drug similarity network (or disease-disease similarity network), which is an $n\times n$ ($m\times m$) real matrix such that the entry $s^{r}_{ij}$ (or $s^{d}_{ij}$) $\in[0,1]$ denotes the degree of similarity of drug $i$ and drug $j$ (or disease $i$ and disease $j$), where $i,j\in\{1,2,\ldots,n\}$ (or $i,j\in\{1,2,\ldots,m\}$). For a drug-disease association network, we use a $n\times m$ binary matrix, denoted by $A$, to represent it, where the entry $a_{ij}=1$ if drug $i$ is associated with disease $j$ and $a_{ij}=0$ if drug $i$ is not associated with disease $j$ or the associativity is yet to be observed, for $i=1,2,\ldots,n$ and $j=1,2,\ldots,m$.
To reduce the impact of data noise, we aggregate only the top-$t$ neighbors for each drug in $S^{r}$ and each disease in $S^{d}$, and then two new matrices $R\in\mathbb{R}^{n\times n}$ and $D\in\mathbb{R}^{m\times m}$ are generated to replace $S^{r}$ and $S^{d}$, respectively, where the entries $r_{ij}$ ($i,j\in\{1,2,\ldots,n\}$) of $R$ and $d_{ij}$ ($i,j\in\{1,2,\ldots,m\}$) of $D$ are defined as follows. For a drug $r_{i}$ (and a disease $d_{i}$), we denote by $N_{t}(r_{i})$ (and $N_{t}(d_{i})$) the set of $t$-nearest neighbors of $r_{i}$ in $S^{r}$ (and $d_{i}$ in $S^{d}$), i.e., the weight of edge $r_{i}x$ (and edge $d_{i}y$) for every $x\in N_{t}(r_{i})$ (and every $y\in N_{t}(d_{i})$) is top-$t$ among all neighbors of $r_{i}$ (and $d_{i}$), and let $\tilde{N_{t}}(r_{i})=N_{t}(r_{i})\cup\{r_{i}\}$ (and $\tilde{N_{t}}(d_{i})=N_{t}(d_{i})\cup\{d_{i}\}$) be the extended $t$-nearest neighbors of $r_{i}$ (and $d_{i}$).
$\begin{matrix}r_{ij}=\begin{cases}1,&\mbox{if }r_{j}\in{\tilde{N_{t}}(r_{i})}\\ 0,&\text{otherwise}\end{cases}\quad&d_{ij}=\begin{cases}1,&\mbox{if }d_{j}\in{\tilde{N_{t}}(d_{i})}\\ 0,&\text{otherwise}\end{cases}\end{matrix}$ (1)